OpenStudy (ineedhelpnow):

an american football field is a rectangle with a perimeter of 1040 feet. the length is 200 feet more than the width. find the width and length of the rectangle field.

OpenStudy (phi):

 perimeter is the sum of the sides: w+L+w+L= 1040 feet this simplifies to 2w+2L= 1040 divide both sides of the equation by 2 (to simplify it even more) to get w+L= 520 also, the length L is 200 feet more than w, or L= 200+w replace L in the equation with 200+w (because L= 200+w) w+200+w= 520 add the w's 2w+200= 520 subtract 200 from each side 2w= 320 divide both sides by 2 to get w. then use w in w+L=520 to find L